OpenStudy (anonymous):

Simplify completely( x^2 + x - 12 / x^2 - x - 20) * (12x^2 -48x -60 / 3x^2 - 24x + 45) Please best answer I will give a metal<3

OpenStudy (anonymous):

factor out each expression first. then simplify. it goes like this.

OpenStudy (anonymous):

you simplify 12x2-48x-60. divide it by 12 since all of them are divisible by 12. you also simplify 3x2-24x+45 by dividing it by 3. you'll get x2-4x-5 and x2-8x+15.

OpenStudy (anonymous):

Thanks you!! :D

OpenStudy (anonymous):

no problem. :)

OpenStudy (anonymous):

@Tabatabachoii You're incorrect. \[\frac{ (x+4)(x-3) }{ (x-5)(x+4) }\times \frac{ 12(x-5)(x+1) }{ 3(x-3)(x-5) }\] \[=\frac{ 12(x+1) }{ 3(x-5) }\] \[=\frac{ 4(x+1) }{ x-5 }\]

OpenStudy (anonymous):

@itsladyrw You must not eliminate the 12 and 3 when you factor out that part of the expression.

OpenStudy (anonymous):

ohw, I forgot to. I admit my mistake. Sorry.

OpenStudy (anonymous):

I assume it when I saw the answer didnt match at my answer key lol, and thanks you @Azteck.